A 1Kg object is tied with a 2m rope and making uniform circular motion on the horizontal surface. the coefficient of friction is 0.2. if the rope can succeed maximum tension of 20N, what is the maximum velocity of the object


1
Expert's answer
2021-01-25T13:56:31-0500

Lets  solve the problem

First of all we write the x and y motion equations

the object travel in a uniform circular motion in a horizontal surface

y - coordinate -> "N = F_{net} + T"

x - coordinate -> "F_{fric} = ma"

"F_{fric} = \mu N = \mu (F_{net} + T) = \mu (ma+T) = ma"

"ma - \mu ma = \mu T \to a = \large\frac{\mu T}{m(1-\mu)} = \frac{v^2}{R}"

"v = \large\sqrt{\frac{\mu TR }{m(1-\mu)}} = \sqrt{\frac{0.2*20*2}{1*(1-0.2)}}" = "3.16 \frac{m}{s}"